# Practical Guidelines for the Selection and Evaluation of NLP Techniques in RE

### v1

@@ Abstract
Natural Language Processing (NLP) is now a cornerstone of requirements automation. One compelling factor behind the growing adoption of NLP in Requirements Engineering (RE) is the prevalent use of natural language (NL) for specifying requirements in industry. NLP techniques are commonly used for automatically classifying requirements, extracting important information, e.g., domain models and glossary terms, and performing quality assurance tasks, such as ambiguity handling and completeness checking. With so many different NLP solution strategies available and the possibility of applying machine learning alongside, it can be challenging to choose the right strategy for a specific RE task and to evaluate the resulting solution in an empirically rigorous manner. This book chapter presents guidelines for the selection of NLP techniques as well as for their evaluation in the context of RE. In particular, we discuss how to choose among different strategies such as traditional NLP, feature-based machine learning, and language-model-based methods. Our ultimate hope for this chapter is to serve as a stepping stone, assisting newcomers to NLP4RE in quickly initiating themselves into the NLP technologies most pertinent to the RE field.
@@ 2 Automation Steps in NLP4RE
The automation process for NL requirements and related textual artifacts can be structured into three sequential steps. These steps, shown in Fig. 1 are: (1) Pre-processing, (2) Analysis, and (3) Post-processing. We outline these steps next.
@@ 2.2 Analysis
The core of the process shown in Fig. 1 is the Analysis step. In NLP4RE, this step typically manifests as one of the following three alternative activities: Classification, Clustering, and Text Generation.
- Classification
involves the assignment of labels or categories to the units of analysis. There are numerous NLP4RE use cases for classification. An example use case would be differentiating between functional (F) and non-functional requirements (NF) [27]. This task can be framed as the assignment of F and NF labels to the units of analysis, which in this context, are typically sentences within a requirements document.
@@
- Clustering
leverages inherent similarities among the units of analysis to organize them into groups or themes. Unlike classification, which requires predetermined labels, clustering emphasizes the intrinsic structure and similarities within the content. The goal is to bring together similar content based on shared features, thereby avoiding the need for explicit predefined categories. This makes clustering particularly useful when dealing with problems where the labels are not well-defined or when exploring content where the underlying patterns might not be immediately evident. For instance, clustering can be used to identify groups of closely related requirements phrases during glossary construction [7]; here there is no predetermined set of classes for the groups of related phrases that will emerge.
- Text Generation
involves the automated creation of human-readable text based on either structured or unstructured inputs to aid the derivation, completion, understanding and communication of requirements. NLP4RE solutions based on text generation are a relatively recent development but are rapidly gaining momentum thanks to advances in generative language models like GPT [9]. Example use cases for text generation include constructing requirements models based on prompts and early-stage descriptions [16], summarizing requirements-related documents [26], and providing predictive assistance for requirements completion [31].
Selecting suitable enabling technique(s) for the Analysis step is crucial. To make this selection easier, we have developed a decision process, shown in Fig. 2. This simple process, which is based on our past experience, aims to facilitate narrowing the options for the Analysis step.
The first and most important criterion in this process is decision node (a), as depicted in Fig 2. This decision concerns whether we have a well-established and pre-existing set of conceptual categories relevant to automation. For instance, consider the task of classifying functional and non-functional requirements. In this task, the categories would be functional and non-functional, and this understanding exists before classification. For another example, consider the task of domain model extraction. Here, all pertinent categories are identified and can be listed, such as class, attribute, association, cardinality constraint, and so on. In contrast, consider a problem like identifying similar and potentially redundant requirements. When presented with a requirements document, predicting the number of equivalence classes (clusters) is impossible. As a result, there is no predefined set of conceptual categories for this problem that can be known beforehand.
@@
When an a-priori-known set of conceptual categories exists, the next question is whether a labelled dataset is available. This is captured by decision node (c) in Fig. 2. When a labelled dataset is lacking, the most common enabling technique is query-based pattern matching. Although pattern matching does not require the explicit creation of labelled data, formulating the queries often entails some form of qualitative analysis, e.g., grounded theory [41]. Alternatively, language models can be employed when labelled data is unavailable. In such a scenario, one has to rely exclusively on the language model’s pre-training, without further specific training for the task at hand. For example, a language model could be asked the following: “Is this requirements statement functional or non-functional? [an individual requirements statement]”.
If labelled data is available, the next factor to consider is the volume of such data, as captured by decision node (d) in Fig. 2. When a considerable amount of labelled data is available, there are two options: applying feature-based learning or utilizing the labelled data to “fine-tune” a language model. Fine-tuning involves adapting a pre-trained model to a specific task through targeted training. In cases of limited labelled data, achieving task adaptation for a language model with minimal examples – a technique known as few-shot learning – is likely to yield better results.
@@
We need to highlight three important aspects related to the process in Fig. 2. First, no individual decision model can encompass the full spectrum of techniques employed in NLP4RE. Our model aims to offer a simplified representation of common scenarios, rather than imposing constraints or promoting a lack of flexibility in technique selection. Second, the process is likely to evolve in the future to stay in step with NLP4RE research. In particular, capitalizing on the interactive capabilities of large language models, there is potential to further elaborate the process by considering prompting strategies and providing additional guidelines. However, due to the scarcity of NLP4RE approaches built on large language models, we have to defer doing so to the future. Finally, when selecting enabling techniques for analysis, the cost and environmental impact must be considered. Most notably, the resource-intensive nature of large language models requires justification, especially when alternatives cannot be dismissed due to compelling reasons such as lack of accuracy.
@@ 2.3 Post-processing
Post-processing – the third step in the process of Fig. 1 – aims to enhance the results of the Analysis step or to adapt these results for human analysts’ better understanding. Post-processing is not needed in all NLP4RE solutions and is thus an optional step.
@@ 3 Enabling Techniques: Overview and Guidelines
In this section, we outline the various enabling techniques shown in Fig. 1 and provide practical guidelines for applying and evaluating them.
@@ 3.1 NLP Pipeline
Figures 3 (a) and (b) respectively exemplify the annotations generated by the constituency parsing and dependency parsing modules for the following requirements sentence: . The annotations generated by the constituency parser for this sentence are: NP (noun phrase), VP (verb phrase), and PP (prepositional phrase). These annotations capture the hierarchical constituents of the sentence. The relationships between the words in the sentence are given by the pairwise links generated through dependency parsing. For instance, from the dependency parse graph of Fig. 3 (b), one can determine that “simulator” functions as the subject (nsubj) for “store,” and that “messages” serves as the object (obj) for this transitive verb. Crucially, both parsing methods require sentence splitting and POS tagging. Figure 3(a) illustrates sentence annotation (S), while both figures include POS tags (DT: determiner, NN: noun, NNS: plural noun, MD: modifier, VB: verb, IN: preposition).
@@
Lemma 1 (NLP Pipeline).
Prioritize pipeline implementations that are known to work well for the language(s) you need to support. Experimentation with different NLP modules is often useful for improving accuracy. If extensive experimentation with the NLP pipeline is not possible or you are limited to a single library, conduct an error analysis on the annotations to identify and address any systemic issues such as recurring errors in sentence detection, POS tags, and parsing.
@@ 3.2 Relevance Measures
Lemma 2 (Relevance Measures).
Combining different relevance measures often results in more accurate analytical outcomes. Relevance measures can be applied at different levels of granularity, ranging from individual tokens (e.g., Levenshtein distance) to entire documents (e.g., TF-IDF). Empirical evaluations of relevance measures can centre around identifying the most effective combination of these measures or benchmarking advanced solutions against relevance measures considered as baseline methods.
@@ 3.3 Embeddings
Lemma 3 (Embeddings).
Experiment with different embedding technologies and dimensionality options to strike a balance between accuracy and the challenges of high-dimensional data. Assess whether the added accuracy of contextual embeddings justifies their higher computational cost, or if non-contextual embeddings suffice for your specific task. Explore domain-specific embeddings when available, as they may be superior at capturing context-specific meanings.
@@ 3.4 Query-based Pattern Matching
Lemma 4 (Query-based Pattern Matching).
(1) When working with NLP annotations, select query language(s) that align with your specific analysis needs. (2) To ensure the reliability of query derivation, carefully document the steps you take to create the queries. This documentation should include the criteria, context, and content used for query design, making it easier for others to reproduce your work. (3) Refrain from using all the text you have for query development. To improve the applicability of your queries, evaluate their performance on unseen (or withheld) text or in different domains.
@@ 3.5 Feature-based Machine Learning
Lemma 5 (Feature-based ML).
Develop explicit procedures for manual data labelling during data collection and document these procedures. Keep track of data-wrangling decisions, as they can have a substantial impact on the outcomes. Approach features as hypotheses and confirm their significance before incorporating them into the final solution. Given the numerous combinations available for hyperparameter tuning during model selection and tuning, you may want to begin with experimentation using the default settings of machine learning algorithms. Ensure there is no data leakage, which means avoiding the inadvertent exposure of parts of test documents during the training and validation processes. The choice of evaluation metrics is of great importance and should be aligned with the specific NLP4RE task at hand.
@@ 3.6 Clustering Algorithms
Lemma 6 (Clustering Algorithms).
Important decisions in clustering include determining a suitable number of clusters (e.g., the Elbow method or Silhouette analysis), and opting between clustering types (partitional, hierarchical, or soft). The choice of clustering algorithm depends on the goal of the analysis and requires domain understanding. Evaluating the effectiveness of clustering is critical, which can be approached internally by assessing the cohesion and separation of clusters or externally by comparing against a ground truth. An external evaluation often poses challenges, as creating a ground truth for clustering is subjective and often requires substantial effort.
@@ 3.7 Large Language Models
Lemma 7 (Large Language Models).
When working with LLMs, explore hyperparameters within resource constraints. Invest in prompt engineering, using suitable prompt patterns and techniques, and empirically compare different alternatives. Take note of the non-deterministic nature of fine-tuning, and present results averaged over multiple fine-tuning runs.
4 Conclusion

### v2

@@ Abstract
[Context and Motivation] Natural Language Processing (NLP) is now a cornerstone of requirements automation. One compelling factor behind the growing adoption of NLP in Requirements Engineering (RE) is the prevalent use of natural language (NL) for specifying requirements in industry. NLP techniques are commonly used for automatically classifying requirements, extracting important information, e.g., domain models and glossary terms, and performing quality assurance tasks, such as ambiguity handling and completeness checking. With so many different NLP solution strategies available and the possibility of applying machine learning alongside, it can be challenging to choose the right strategy for a specific RE task and to evaluate the resulting solution in an empirically rigorous manner. [Content] In this chapter, we present guidelines for the selection of NLP techniques as well as for their evaluation in the context of RE. In particular, we discuss how to choose among different strategies such as traditional NLP, feature-based machine learning, and language-model-based methods. [Contribution] Our ultimate hope for this chapter is to serve as a stepping stone, assisting newcomers to NLP4RE in quickly initiating themselves into the NLP technologies most pertinent to the RE field.
@@ 2 Automation Steps in NLP4RE
The automation process for NL requirements and related textual artifacts can be structured into three sequential steps. These steps, shown in Figure 1 are: (1) Pre-processing, (2) Analysis, and (3) Post-processing. We outline these steps next.
@@ 2.2 Analysis
The core of the process shown in Figure 1 is the Analysis step. In NLP4RE, this step typically manifests as one of the following three alternative activities: Classification, Clustering, and Text Generation.
2.2.1 Analysis Activities
Classification involves the assignment of labels or categories to the units of analysis. There are numerous NLP4RE use cases for classification. An example use case would be differentiating between functional (F) and non-functional requirements (NF) [27]. This task can be framed as the assignment of F and NF labels to the units of analysis, which in this context, are typically sentences within a requirements document.
@@ 2.2.1 Analysis Activities
Clustering leverages inherent similarities among the units of analysis to organize them into groups or themes. Unlike classification, which requires predetermined labels, clustering emphasizes the intrinsic structure and similarities within the content. The goal is to bring together similar content based on shared features, thereby avoiding the need for explicit predefined categories. This makes clustering particularly useful when dealing with problems where the labels are not well-defined or when exploring content where the underlying patterns might not be immediately evident. For instance, clustering can be used to identify groups of closely related requirements phrases during glossary construction [7]; here there is no predetermined set of classes for the groups of related phrases that will emerge.
Text Generation involves the automated creation of human-readable text based on either structured or unstructured inputs to aid the derivation, completion, understanding and communication of requirements. NLP4RE solutions based on text generation are a relatively recent development but are rapidly gaining momentum thanks to advances in generative language models like GPT [9]. Example use cases for text generation include constructing requirements models based on prompts and early-stage descriptions [16], summarizing requirements-related documents [26], and providing predictive assistance for requirements completion [31].
2.2.2 Technique Selection
Selecting suitable enabling technique(s) for the Analysis step is crucial. To make this selection easier, we have developed a decision process, shown in Figure 2. This simple process, which is based on our past experience, aims to facilitate narrowing the options for the Analysis step.
The first and most important criterion in this process is decision node (a), as depicted in Figure 2. This decision concerns whether we have a well-established and pre-existing set of conceptual categories relevant to automation. For instance, consider the task of classifying functional and non-functional requirements. In this task, the categories would be functional and non-functional, and this understanding exists before classification. For another example, consider the task of domain model extraction. Here, all pertinent categories are identified and can be listed, such as class, attribute, association, cardinality constraint, and so on. In contrast, consider a problem like identifying similar and potentially redundant requirements. When presented with a requirements document, predicting the number of equivalence classes (clusters) is impossible. As a result, there is no predefined set of conceptual categories for this problem that can be known beforehand.
@@ 2.2.2 Technique Selection
When an a-priori-known set of conceptual categories exists, the next question is whether a labelled dataset is available. This is captured by decision node (c) in Figure 2. When a labelled dataset is lacking, the most common enabling technique is query-based pattern matching. Although pattern matching does not require the explicit creation of labelled data, formulating the queries often entails some form of qualitative analysis, e.g., grounded theory [41]. Alternatively, language models can be employed when labelled data is unavailable. In such a scenario, one has to rely exclusively on the language model’s pre-training, without further specific training for the task at hand. For example, a language model could be asked the following: “Is this requirements statement functional or non-functional? [an individual requirements statement]”.
If labelled data is available, the next factor to consider is the volume of such data, as captured by decision node (d) in Figure 2. When a considerable amount of labelled data is available, there are two options: applying feature-based learning or utilizing the labelled data to “fine-tune” a language model. Fine-tuning involves adapting a pre-trained model to a specific task through targeted training. In cases of limited labelled data, achieving task adaptation for a language model with minimal examples – a technique known as few-shot learning – is likely to yield better results.
@@
We need to highlight three important aspects related to the process in Figure 2. First, no individual decision model can encompass the full spectrum of techniques employed in NLP4RE. Our model aims to offer a simplified representation of common scenarios, rather than imposing constraints or promoting a lack of flexibility in technique selection. Second, the process is likely to evolve in the future to stay in step with NLP4RE research. In particular, capitalizing on the interactive capabilities of large language models, there is potential to further elaborate the process by considering prompting strategies and providing additional guidelines. However, due to the scarcity of NLP4RE approaches built on large language models, we have to defer doing so to the future. Finally, when selecting enabling techniques for analysis, the cost and environmental impact must be considered. Most notably, the resource-intensive nature of large language models requires justification, especially when alternatives cannot be dismissed due to compelling reasons such as lack of accuracy.
@@ 2.3 Post-processing
Post-processing – the third step in the process of Figure 1 – aims to enhance the results of the Analysis step or to adapt these results for human analysts’ better understanding. Post-processing is not needed in all NLP4RE solutions and is thus an optional step.
@@ 3 Enabling Techniques: Overview and Guidelines
In this section, we outline the various enabling techniques shown in Figure 1 and provide practical guidelines for applying and evaluating them.
@@ 3.1 NLP Pipeline
Figures 3 (a) and (b) respectively exemplify the annotations generated by the constituency parsing and dependency parsing modules for the following requirements sentence: . The annotations generated by the constituency parser for this sentence are: NP (noun phrase), VP (verb phrase), and PP (prepositional phrase). These annotations capture the hierarchical constituents of the sentence. The relationships between the words in the sentence are given by the pairwise links generated through dependency parsing. For instance, from the dependency parse graph of Figure 3 (b), one can determine that “simulator” functions as the subject (nsubj) for “store,” and that “messages” serves as the object (obj) for this transitive verb. Crucially, both parsing methods require sentence splitting and POS tagging. Figure 3(a) illustrates sentence annotation (S), while both figures include POS tags (DT: determiner, NN: noun, NNS: plural noun, MD: modifier, VB: verb, IN: preposition).
@@ 3.7 Large Language Models
4 Summary and Conclusion